# Evacuation Planning on Time-Expanded Networks with Integrated Wildfire Information

### v1

@@ 2.3 Our Contribution
Once constructed, we solve a maximum flow problem on the WTEN. This can be formulated as a linear program, (see WTEN LP), and for a convenient wording, we say that we solve (WTEN LP) even though in practice we are using a more efficient combinatorial network flow algorithm. While there do exist combinatorial algorithms for solving problems on the original dynamic network, their efficiency relies on certain restrictions. These restrictions could force the oversimplification of large-scale evacuation models. For example, simply adding or removing arcs cannot capture changes in arc capacity or travel time which are needed to represent changing road conditions during an evacuation. This serves as additional motivation for utilizing time-expanded networks in our setting.
In Section 4, we provide a high-level explanation of each step in our algorithm. First, we describe the pre-processing of public network data used to construct the WTEN. This includes using the Python packages OSMNX, [6], and NetworkX. We start by taking data from OpenStreetMaps [44] and create a network object. Node and arc information such as road capacities, supply and demand values (which represent the population we need to evacuate and the capacity of evacuation locations), and longitudinal/latitudinal coordinates are added to the network. Next, a sample fire is created and used to construct a WTEN; see Section 3. We find an optimal solution to (WTEN LP) using the shortest augmenting path algorithm. This solution serves as an initial evacuation plan. We confirm the overall running time and prove correctness, in particular correct termination, of the approach.
@@ 3 Integration of Hazard Information in Time-Expanded Networks
Our modeling goal is the integration of dynamic hazard information, in particular wildfire information, into a time-expanded network. In Section 2.1, we saw that the key steps when converting a dynamic network to a time-expanded network are the discretization of time and the copying of nodes and arcs. We note that the number of nodes in a time-expanded network scales directly with the granularity of the discretized time horizon. In our proof of concept computations in Section 5, we typically observe time horizons of 20-50 minutes to complete a regional evacuation, and use a minute as a unit of time to balance model detail and efficiency of solution. Travel times along arcs are rounded up to the nearest integer.
Depending on the spread of a wildfire, nodes in the time-expanded network may not be visited at certain time instances, and arcs may have reduced capacity or may not be used. Therefore, to accurately model hazard information in the time-expanded network, we cannot simply copy our node and arc sets for each time step. Instead, we must incorporate additional information, i.e., the time at which each node and arc is impacted by the hazard. There exist various prediction models for different hazards, such as wildfires and floods. Our algorithm can incorporate any model that, for each time instance $t$ in the discretized time horizon, provides the area or areas that are impacted by the hazard; we denote this set $F(t)$. The necessary notation for the underlying TEN has been introduced in Section 2. Here we add notation regarding the fire prediction model used in conjunction with TEN, to yield a new network WTEN, the wildfire time-expanded network.
@@
$N_{F(t)}:=\left\{i\in N:\left({\text{lat}}_{i},{\text{long}}_{i}\right)\in F(t)\right\},$
@@
$N_{W}:=\left\{i(t)\in N_{T}:i\in N\setminus N_{F(t)},\ t\in[T]_{0}\right\}.$
@@
$t_{\text{max}}(j):=\max\{t\in[T]_{0}:d_{j}\notin N_{F(t)}\}$
$S_{W}:=\{s_{i}(0)\in N_{W}:s_{i}(0)\in S_{T}\setminus N_{F(0)}\}\;\;\text{and}$
$D_{W}:=\{d_{j}(t_{\text{max}}(j))\in N_{W}:d_{j}\in D\}$
@@
$A^{M}_{W}$ :=$:=\{(i(t),j(t^{\prime})\in A^{M}:(i,j)\in A,\ i\in N\setminus N_{F(t)},\ j\in N\setminus N_{F(t^{\prime})},\ t^{\prime}=t+\lambda_{ij}(t)\leq T,\ t\in[T-1]_{0}\}$
$A^{H}_{W}$ :=$:=\{(i(t),i(t+1)\in A^{H}:i\in(S_{W}\cup D_{W})\setminus(N_{F(t)}\cup N_{F(t+1)}),\ t\in[T-1]_{0}\},$
@@
Next, we adjust arc capacities for movement arcs remaining in the WTEN. Specifically, we assume that the distance from a road to the fire directly affects how many people are able to safely travel along the road. One could model this by increasing travel times $\lambda_{ij}(t)$ or reducing capacities $u_{ij}(t)$ over time. We elect to model it by reducing the upper capacities $u_{ij}(t)$ of movement arcs close to the fire, and keep the travel times $\lambda_{ij}(t)$ along an arc the same. This implies that the updated $A^{M}_{W}$ remains fixed, and only capacities need to be adjusted. In our fire prediction model, we also assume the fire is growing at a rate of one meter per minute [2]. This allows us to directly compare travel times to the distance to the fire without needing to convert units. We define the parameter $p_{ij}(t)$ as the percentage of available capacity of arc $(i,j)\in A$ at time $t$:
$p_{ij}(t)=\begin{cases}1&f_{ij}(t)\geq\lambda_{ij}(t)\\ \frac{f_{ij}(t)}{\lambda_{ij}(t)}&f_{ij}(t)<\lambda_{ij}(t),\end{cases}$
@@
$u^{W}_{i(t)j(t^{\prime})}=u_{ij}(t)=\lfloor u_{ij}(0)\cdot p_{ij}(t)\rfloor,$
where $i(t),j(t^{\prime})\in N_{W}$ and $t^{\prime}=t+\lambda_{ij}(t)$ are some copy of $i,j\in N$ in the WTEN such that $(i(t),j(t^{\prime}))\in A^{M}_{W}$. In our model, holdover arcs are treated as having the same capacity for all time instances, i.e., $a_{i}(0)=a_{i}(t)=a_{i}$ for all $1\leq t\leq T$ and for all $i\in(S_{W}\cup D_{W})$. However, it would be simple to adjust these capacities based on a distance $f_{ii}(t)$ between $i\in N$ and $F(t)$ if desired.
@@
$A^{(S_{W}\cup D_{W})}_{W}:=\left\{\left(\hat{s},j\right):j\in S_{W}\right\}\cup\{(i,\hat{d}):i\in D_{W}\}.$
@@
$N_{W}:=\left\{i(t)\in N_{T}:i\in N\setminus N_{F(t)},\ t\in[T]_{0}\right\}\cup\{\hat{s},\hat{d}\}\text{ and }A_{W}:=A^{M}_{W}\cup A^{H}_{W}\cup A^{(S_{W}\cup D_{W})}_{W}$
@@
$A^{M}_{W}$ :=$:=\{(i(t),j(t^{\prime}))\in A^{M}:(i,j)\in A,\ i\in N\setminus N_{F(t)},\ j\in N\setminus N_{F(t^{\prime})},\ t^{\prime}=t+\lambda_{ij}(t)\leq T,t\in[T-1]_{0}\}$
@@
$u^{W}_{i(t)j(t^{\prime})}$ $:=\lfloor u_{ij}(0)p_{ij}(t)\rfloor$ for $\text{for }i,j\in A,\ t^{\prime}=\lambda_{ij}(t)+t,\ t\in[T-1]_{0}$
$u^{W}_{i(t)i(t+1)}$ $:=a_{i}(0)$ for $\text{for }i\in(S\cup D),\ t\in[T-1]_{0}$
$u^{W}_{\hat{s}j^{\prime}}$ $:=a_{s_{j}}(0)$ for $\text{for }j^{\prime}\in S_{W},\ s_{j}\in S,\ j^{\prime}=s_{j}(0)$
$u^{W}_{i^{\prime}\hat{d}}$ $:=a_{d_{i}}(0)$ for $\text{for }i^{\prime}\in D_{W},\ d_{i}\in D,\ i^{\prime}=d_{i}(t_{\text{max}}(i)).$
@@
max ($\begin{array}[]{rcrcrcl}\text{max }&&\displaystyle{\sum_{(i,\hat{d})\in A_{W}}}x_{i{\hat{d}}}&&&&\\ \text{s. t. }&&\displaystyle{\sum_{(j,i)\in A_{W}}}x_{ji}-\displaystyle{\sum_{(i,j)\in A_{W}}}x_{ij}&=&0&&\forall i\in N_{W}\setminus\{\hat{s},\hat{d}\}\\ &&0\leq x_{ij}&\leq&u^{W}_{ij}&&\forall(i,j)\in A_{W}.\end{array}$ (WTEN LP)
@@ 3.1 Wildfire Modeling
Traditional wildfire models can be computationally expensive or require expert knowledge to use. Moreover, with the availability of more and better data, wildfire models have grown in sophistication, detail, and accuracy. Today, many models take into account weather conditions, amount and types of burnable materials, and even flame height and length [45, 48].
In an effort to broaden usability, the code for our algorithm allows for the incorporation of any wildfire model via a GIS shapefile output that represents where the fire is, i.e., a fire set $F(t)$, at a given time, $t$. That is, any geometry or shapefile that is compatible with the Python Shapely package can be integrated with our algorithm. A detailed, state-of-the-art fire model can be used to generate the shapefile for an initial plan, where computational speed is not a bottleneck. Those plans can also be pre-computed and used as a template. However, during a developing fire and evacuation, computational speed is vital to enable a rapid reaction to a change in fire spread. In such a situation, one would need to use a simple fire model to avoid a bottleneck.
@@
$F(t):=\left\{(x,y)\in\mathbb{R}^{2}:(x,y)\in\bigcup_{k=1}^{\ell}C_{k}\right\},$
@@ 4 The Algorithm
Input location information Construct original network Compute initial time horizon $T$ Model fire growth Construct WTEN Solve (WTEN LP) Everyone evacuated? $T=T+1$ $T>T_{\text{max}}$ Return maximum flow solution for time horizon $T$ $T=T_{\text{sol}}$ nonoyesyes
@@ 4.1 Data and Pre-processing
OpenStreetMaps is an open source GIS database of geographic and road information [44]. It can produce road data for geographic locations based on various inputs such as bounding boxes, a select distance from a center point, or a city/county name. For more information regarding OpenStreetMaps data and license and OMSNX package we refer the reader to [6, 44]. Since OpenStreetMaps is compiled by the public, road data may not be available for all input locations, however, a major benefit of open source data is that it is updated regularly and therefore can lead to greater accuracy.
@@
The network object is updated to include the parameters for a dynamic network; see Section 2. This information can be appended to the network object using built-in OMSNX functions. Travel times can be added directly from OpenStreetMaps (which are rounded to integers during the TEN creation), while the computation of road capacities requires additional steps. First, the speed for each road is added to the network. The capacity for each road is then computed using the Moore Method [40], a simple calculation using empirical data on safe vehicle distance; it is rounded down and attached to the corresponding arc. Finally, each node may be given a supply or demand value, determined by user input. A supply value identifies a source, where the corresponding number of people start their evacuation. A demand value identifies a sink; a destination location with the corresponding capacity.
Next, we establish an initial time horizon $T$. We are looking for a minimal $T$ that will allow for a complete evacuation. A lower bound on this $T$ is the longest of the shortest paths from any source to its closest sink in the original network (based on travel time). There are $|S|\cdot|D|\leq\left(\frac{|N|}{2}\right)^{2}$ such paths to compute, each of them requiring effort $O((|A|+|N|)\log|N|)$ via Dijkstra’s Algorithm; see for example [1]. We find this value using a built-in NetworkX function which computes the shortest path using original travel times between any two nodes in the dynamic network.

### v2

@@ 2.3 Our Contribution
Once constructed, we solve a maximum flow problem on the WTEN. This can be formulated as a linear program, (see LABEL:WTEN_LP), and for a convenient wording, we say that we solve (LABEL:WTEN_LP) even though in practice we are using a more efficient combinatorial network flow algorithm. While there do exist combinatorial algorithms for solving problems on the original dynamic network, their efficiency relies on certain restrictions. These restrictions could force the oversimplification of large-scale evacuation models. For example, simply adding or removing arcs cannot capture changes in arc capacity or travel time which are needed to represent changing road conditions during an evacuation. This serves as additional motivation for utilizing time-expanded networks in our setting.
In Section 4, we provide a high-level explanation of each step in our algorithm. First, we describe the pre-processing of public network data used to construct the WTEN. This includes using the Python packages OSMNX, (Boeing, 2017), and NetworkX. We start by taking data from OpenStreetMaps (OpenStreetMap contributors, 2024) and create a network object. Node and arc information such as road capacities, supply and demand values (which represent the population we need to evacuate and the capacity of evacuation locations), and longitudinal/latitudinal coordinates are added to the network. Next, a sample fire is created and used to construct a WTEN; see Section 3. We find an optimal solution to (LABEL:WTEN_LP) using the shortest augmenting path algorithm. This solution serves as an initial evacuation plan. We confirm the overall running time and prove correctness, in particular correct termination, of the approach.
@@ 3 Integration of Hazard Information in Time-Expanded Networks
Our modeling goal is the integration of dynamic hazard information, in particular wildfire information, into a time-expanded network. More specifically, we aim to construct a modular means of integrating information on any hazard into a time-expanded network. As a proof of concept, we focus on modeling wildfire evacuations using a built-in basic fire prediction model. In doing so, we exhibit the type and format of information required by our algorithm that real-world prediction models would need to provide in order to be incorporated.
In Section 2.1, we saw that the key steps when converting a dynamic network to a time-expanded network are the discretization of time and the copying of nodes and arcs. We note that the number of nodes in a time-expanded network scales directly with the granularity of the discretized time horizon. In our proof of concept computations in Section 5, we typically observe time horizons of 20-50 minutes to complete a regional evacuation, and use a minute as a unit of time to balance model detail and efficiency of solution. Travel times along arcs are rounded up to the nearest integer.
Depending on the spread of a wildfire, nodes in the time-expanded network may not be visited at certain time instances, and arcs may have reduced capacity or may not be used. Therefore, to accurately model hazard information in the time-expanded network, we cannot simply copy our node and arc sets for each time step. Instead, we must incorporate additional information, i.e., the time at which each node and arc is impacted by the hazard. There exist various prediction models for different hazards, such as wildfires and floods. Our algorithm can incorporate any model that, for each time instance $t$ in the discretized time horizon, provides the area or areas that are impacted by the hazard; we denote this set $F(t)$. The necessary notation for the underlying TEN has been introduced in Section 2. Here we add notation regarding the fire prediction model used in conjunction with TEN, to yield a new network WTEN, the wildfire time-expanded network. Note that we use this name for ease, in reality, “wildfire” may be replaced with any appropriate hazard.
@@
$N_{F(t)}:=\left\{i\in N:\left({\text{lat}}_{i},{\text{long}}_{i}\right)\in F(t)\right\},$ (17)
@@
$N_{W}:=\left\{i(t)\in N_{T}:i\in N\setminus N_{F(t)},\ t\in[T]_{0}\right\}.$ (18)
@@
$t_{\text{max}}(j):=\max\{t\in[T]_{0}:d_{j}\notin N_{F(t)}\}$ (19)
$S_{W}:=\{s_{i}(0)\in N_{W}:s_{i}(0)\in S_{T}\setminus N_{F(0)}\}\;\;\text{and}$ (20)
$D_{W}:=\{d_{j}(t_{\text{max}}(j))\in N_{W}:d_{j}\in D\}$ (21)
@@
$A^{M}_{W}$ :=$:=\{(i(t),j(t^{\prime}))\in A^{M}:(i,j)\in A,\ i\in N\setminus N_{F(t)},\ j\in N\setminus N_{F(t^{\prime})},$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad t^{\prime}=t+\lambda_{ij}(t)\leq T,\ t\in[T-1]_{0}\}$ (22)
$A^{H}_{W}$ :=$:=\{(i(t),i(t+1))\in A^{H}:i\in(S_{W}\cup D_{W})\setminus(N_{F(t)}\cup N_{F(t+1)}),\ t\in[T-1]_{0}\},$ (23)
@@
Next, we adjust arc capacities for the movement arcs remaining in the WTEN. Specifically, we assume that the distance from a road to the hazard directly affects how many people are able to safely travel along the road in a given time window. This is intuitive, for example, in a fire, smoke may reduce visibility on the road, leading to drivers slowing down. The impact can be interpreted in two ways, either of which could be directly incorporated into our model: drivers take longer to traverse that stretch of road; or fewer people are able to make it from the start to the end of the road during the same travel time. One can model this effect by increasing travel times $\lambda_{ij}(t)$ or reducing capacities $u_{ij}(t)$ over time, respectively. We elect to model it by reducing the upper capacities $u_{ij}(t)$ of movement arcs close to the fire, and keep the travel times $\lambda_{ij}(t)$ along an arc the same. This implies that the updated $A^{M}_{W}$ remains fixed, and only capacities need to be adjusted. In our fire prediction model, we also assume the fire is growing at a rate of one meter per minute (Albini, 1976). This allows us to directly compare travel times to the distance to the fire without needing to convert units. We define the parameter $p_{ij}(t)$ as the percentage of available capacity of arc $(i,j)\in A$ at time $t$:
$p_{ij}(t)=\begin{cases}1&f_{ij}(t)\geq\lambda_{ij}(t)\\ f_{ij}(t)/\lambda_{ij}(t)&f_{ij}(t)<\lambda_{ij}(t),\end{cases}$ (24)
@@
$u^{W}_{i(t)j(t^{\prime})}=u_{ij}(t)=\lfloor u_{ij}(0)\cdot p_{ij}(t)\rfloor,$ (25)
where $i(t),j(t^{\prime})\in N_{W}$ and $t^{\prime}=t+\lambda_{ij}(t)$ are some copy of $i,j\in N$ in the WTEN such that $(i(t),j(t^{\prime}))\in A^{M}_{W}$. In our model, holdover arcs are treated as having the same capacity for all time instances, i.e., $a_{i}(0)=a_{i}(t)=a_{i}$ for all $1\leq t\leq T$ and for all $i\in(S_{W}\cup D_{W})$. However, it would be simple to adjust these capacities based on the distance $f_{ii}(t)$, i.e, the distance between $i\in N$ and $F(t)$, if desired. For example, those source nodes closest to the fire could have reduced holdover capacity between earlier time instances to force evacuees to leave quickly. This type of capacity reduction can also replicate some ideas of a stage-based evacuation in which specific nodes, such as those with the largest at-risk populations, begin evacuating early.
@@
$A^{(S_{W}\cup D_{W})}_{W}:=\left\{\left(\hat{s},j\right):j\in S_{W}\right\}\cup\{(i,\hat{d}):i\in D_{W}\}.$ (26)
@@
$N_{W}:=\left\{i(t)\in N_{T}:i\in N\setminus N_{F(t)},\ t\in[T]_{0}\right\}\cup\{\hat{s},\hat{d}\}\text{ and }A_{W}:=A^{M}_{W}\cup A^{H}_{W}\cup A^{(S_{W}\cup D_{W})}_{W}$ (27)
@@
$A^{M}_{W}$ :=$:=\{(i(t),j(t^{\prime}))\in A^{M}:(i,j)\in A,\ i\in N\setminus N_{F(t)},\ j\in N\setminus N_{F(t^{\prime})},$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad t^{\prime}=t+\lambda_{ij}(t)\leq T,t\in[T-1]_{0}\}$
@@
$u^{W}_{i(t)j(t^{\prime})}$ $:=\lfloor u_{ij}(0)p_{ij}(t)\rfloor$ for $\text{for }i,j\in A,\ t^{\prime}=\lambda_{ij}(t)+t,\ t\in[T-1]_{0}$ (28)
$u^{W}_{i(t)i(t+1)}$ $:=a_{i}(0)$ for $\text{for }i\in(S\cup D),\ t\in[T-1]_{0}$ (29)
$u^{W}_{\hat{s}j^{\prime}}$ $:=a_{s_{j}}(0)$ for $\text{for }j^{\prime}\in S_{W},\ s_{j}\in S,\ j^{\prime}=s_{j}(0)$ (30)
$u^{W}_{i^{\prime}\hat{d}}$ $:=a_{d_{i}}(0)$ for $\text{for }i^{\prime}\in D_{W},\ d_{i}\in D,\ i^{\prime}=d_{i}(t_{\text{max}}(i)).$ (31)
@@
max $\displaystyle{\sum_{(i,\hat{d})\in A_{W}}}x_{i{\hat{d}}}$ (32)
s. t. $\displaystyle{\sum_{(j,i)\in A_{W}}}x_{ji}-\displaystyle{\sum_{(i,j)\in A_{W}}}x_{ij}=0$ $\forall i\in N_{W}\setminus\{\hat{s},\hat{d}\}$ (33)
$0\leq x_{ij}\leq u^{W}_{ij}$ $\forall(i,j)\in A_{W}.$ (34)
@@ 3.1 Wildfire Modeling
Traditional wildfire models can be computationally expensive or require expert knowledge to use. Moreover, with the availability of more and better data, wildfire models have grown in sophistication, detail, and accuracy. Today, many models take into account weather conditions, amount and types of burnable materials, and even flame height and length (Or et al., 2023; Papadopoulos and Pavlidou, 2011). These complex models may not be readily available for all levels of evacuation and response teams, especially those working on a local scale. To accommodate this potential barrier, we provide a routine in our repository allowing the user to model hazard (in particular, fire) growth and impact area over time. This enables anyone with access to our code to obtain clear, implementable evacuation plans based on approximate hazard development, while also allowing experts to easily integrate information obtained from sophisticated wildfire models should they have access to such data.
More precisely, the code for our algorithm allows for the incorporation of any wildfire model via a GIS shapefile output that represents where the fire is, i.e., a fire set $F(t)$, at a given time, $t$. That is, any geometry or shapefile that is compatible with the Python Shapely package can be integrated with our algorithm. A detailed, state-of-the-art fire model can be used to generate the shapefile for an initial plan, where computational speed is not a bottleneck. Those plans can also be pre-computed and used as a template. However, during a developing fire and evacuation, computational speed is vital to enable a rapid reaction to a change in fire spread. In such a situation, one would need to use a simple fire model to avoid a bottleneck.
@@
$F(t):=\left\{(x,y)\in\mathbb{R}^{2}:(x,y)\in\bigcup_{k=1}^{\ell}C_{k}\right\},$ (35)
@@ 4 The Algorithm
Input location information Construct original network Compute initial time horizon $T$ Model fire growth Construct WTEN Solve (LABEL:WTEN_LP) Everyone evacuated? $T=T+1$ $T>T_{\text{max}}$ Return maximum flow solution for time horizon $T$ $T=T_{\text{sol}}$ nonoyesyes
@@ 4.1 Data and Pre-processing
OpenStreetMaps is an open source GIS database of geographic and road information (OpenStreetMap contributors, 2024). It can produce road data for geographic locations based on various inputs such as bounding boxes, a select distance from a center point, or a city/county name. For more information regarding OpenStreetMaps data and license and OMSNX package we refer the reader to (Boeing, 2017; OpenStreetMap contributors, 2024). Since OpenStreetMaps is compiled by the public, road data may not be available for all input locations, however, a major benefit of open source data is that it is updated regularly and therefore can lead to greater accuracy.
@@
The network object is updated to include the parameters for a dynamic network; see Section 2. This information can be appended to the network object using built-in OMSNX functions. Travel times can be added directly from OpenStreetMaps (which are rounded to integers during the TEN creation), while the computation of road capacities requires additional steps. First, the speed for each road is added to the network. The capacity for each road is then computed using the Moore Method (Moore et al., 2013), a simple calculation using empirical data on safe vehicle distance; it is rounded down and attached to the corresponding arc. Finally, each node may be given a supply or demand value, determined by user input. A supply value identifies a source, where the corresponding number of people start their evacuation. A demand value identifies a sink; a destination location with the corresponding capacity.
Next, we establish an initial time horizon $T$. We are looking for a minimal $T$ that will allow for a complete evacuation. A lower bound on this $T$ is the longest of the shortest paths from any source to its closest sink in the original network (based on travel time). There are $|S|\cdot|D|\leq\left(|N|/2\right)^{2}$ such paths to compute, each of them requiring effort $O((|A|+|N|)\log|N|)$ via Dijkstra’s Algorithm; see for example (Ahuja et al., 1993). We find this value using a built-in NetworkX function which computes the shortest path using original travel times between any two nodes in the dynamic network.